Write the slope-intercept form of the equation of the line that passes through the point (-5,4) and has a slope of -1.
Thank you so much!!!!!
God Bless and Happy New Year!!!!!

Answer by rfer(16322)

y=mx+b
4=-1*-5+b
4=5+b
b=-1
------------------
y=mx+b
y=-1*-5-1
y=4
-------------
y=-x-1